# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v1

@@ 5 Performance Analysis of RC-S
In a multi-robot system, the mutual influence among robots often leads to suboptimal utilization of roadway resources. In an ideal scenario, without any robot making turns and merely traveling from a loading station to a designated station, the flow of robots would be uniformly distributed across each link, maximizing network capacity. However, under such circumstances, it becomes unfeasible for a single loading station to serve multiple outlets—a common requirement in sortation scenarios, especially those without pre-sorting. Therefore, we depict the network capacity reduction mechanism by analyzing the characteristics of RC-S, and investigate how the number of operational loading stations and the total number of robots affect the traffic.
@@
Under the logic of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
@@
The speed of each VP could be obtained by expression (4)-(6) with the parameters of RC-S, while the number of occupied VPs is influenced by two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs. Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
@@
where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network. According to the phase analysis in section 4.3, the RC-S scheme stipulates that half of the aisle cells are occupied by VPs at any given time. Therefore, we obtain:
$n_{VP}=\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2}$ (14)
@@
In the operational context, the relationship between a VP and its respective entrance mirrors the relationship between a conveyor belt and its driving electric motor. A VP becomes eligible for occupation only when its associated loading station, situated at the aisle’s entrance, is operational—indicative of the corresponding worker being actively engaged in tasks. The quantity of such active VPs is a critical determinant of the maximum capacity for robot throughput. Take, for instance, a scenario where 8 workers are stationed within an RSS featuring a $6+6$ aisle network. To ensure even distribution of traffic flow and mitigate congestion at the aisle network’s periphery, these workers are strategically positioned along the four edges, as depicted in Figure 9. For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of active workers to the network’s maximum operational capacity in discussions that follow.
@@
Figure 9: The impact of number of workers on the VPs that can be occupied
The top 4 horizontal aisles and left 4 vertical aisles are served by workers, and thus robots could be loaded with parcels at the corresponding entrances. We merge the areas covered by these aisles and obtain the blue polygon. The ratio of area, namely the workforce factor $\kappa$, could be obtained by:
@@
We now examine the path reservation strategy within the RC-S framework. In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation method exhibits a degree of ”inflexibility” because it obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time, thereby diminishing the network’s traffic capacity. To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the proportion of network throughput decrease relative to its potential maximum.
Figure 10: The ”pile up” process of spatiotemporal occupation under RC.
Figure 11: The loss of capacity in different network.
@@
$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2},\ n_{r}\}$ (18)
@@
We now derive the average travel distance of a sorting task, as the remaining part of the throughput estimation formula. Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance. Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns better with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions. Sorting demand are uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery. However, from the diagram, it can be observed that to ensure the outlets in area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$. We will next present the rationale for their calculation and the formulas. For a detailed derivation process, please refer to the appendix B.
@@
[…]
@@
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$. Denote the value as $n_{l}^{(0)}$. Then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 1 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B. Proposition 1 reveals that under the RC-S robot control scheme, the system’s maximum operational efficiency has a critical point, and this critical point is not achieved when all loading stations are open. This may seem counter-intuitive, but remember that we have assumptions of workload balancing and uniformly distributed sorting demands. When any short-sighted dispatch strategy is employed at the loading stations, robots dispatched from the corners will negatively impact the overall throughput of the entire system.
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
Tables 3 and 4 report the system throughput, average service time and average run-time. The comparative results indicate that the runtime of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles.
Table 3: Results on 12+12 aisles directed map
@@
40 8.81 4.72 11.70 14.17 5.26 5.99
80 14.94 9.03 12.90 14.57 6.32 16.85
120 16.99 12.40 13.81 15.24 7.52 35.37
160 17.58 14.32 14.15 15.79 8.07 55.31
200 17.81 14.94 14.32 16.03 8.35 63.36
Table 4: Results on 20+20 aisles directed map
@@
50 7.41 3.94 19.34 21.75 14.38 11.61
100 14.12 7.79 19.92 21.95 14.96 29.94
200 21.04 15.07 20.99 22.54 19.32 99.61
300 22.34 21.18 21.48 23.33 21.30 226.76
400 22.91 24.37 21.97 23.92 22.08 363.45
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail. We conduct 20 rounds of replication for each scenario. In each experiment, we set a warm-up period of one hour and then record the system performance for three hours of continuous operation.
@@
Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste. Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$. Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We notice that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation. Consequently, it incurs a certain efficiency loss, resulting in a decrease in the throughput, as demonstrated in Proposition 1.
@@ 7 Layout Design Optimization
The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
@@
$C_{o}=P_{w}\cdot[\theta\cdot n_{w}^{H}+(1-\theta)\cdot n_{w}^{L}]+P_{r}\cdot[\theta\cdot n_{r}^{H}+(1-\theta)\cdot n_{r}^{L}]$ (31)
Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $M_{w}$ and $M_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot. We assume that there are two typical values of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with $\theta$ denoting the ratio of peak seasons.
The objective of the problem is as follows:
$C_{d}({T}^{L},{T}^{H},\theta,N_{o})=C_{f}+C_{o}$ (32)
@@
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{H},n_{r}^{H})\geq{T}^{H}$ (33)
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{L},n_{r}^{L})\geq{T}^{L}$ (34)
@@
$n_{w}^{H}\leq n_{l}$ (37)
$n_{w}^{L}\leq n_{l}$ (38)
$n_{h}=2k_{h}\quad\forall k_{h}\in\mathcal{Z}_{+}$ (39)
$n_{v}=2k_{v}\quad\forall k_{v}\in\mathcal{Z}_{+}$ (40)
$n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\in\mathcal{Z}_{+}^{(6)}$ (41)
where constraints (33) and (34) ensure that the solution meets predicted sorting requirements. Constraint (35) restricts the lower bound of the site size to accommodate an adequate number of outlets. Constraint (36) represents the maximum number of loading stations accommodated in the aisle network under RC-S. Constraints (37) and (38) limit the maximum number of workers to be stationed. By the design of the system in section 3, we add constraints (39)-(40) to ensure the number of aisles even. The remaining constraints include the expressions to calculate the cost of each component and the estimation formula of system efficiency.
@@
$n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (42)
$n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (43)

### v2

@@ 5 Performance Analysis of RC-S
This section elaborates on the network capacity by analyzing the characteristics of RC-S and investigates how the system configurations affect system throughput. A throughput estimation formula is proposed, serving as a constraint for the planning model in section 7.
@@
Under the mechanism of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
@@
The speed of each VP could be obtained by expression (4)-(5) with the parameters of RC-S. For the number of occupied VPs, an intuitive idea is that it is related to the number of workers engaged in loading parcels in the loading zone and the spatial utilization under RC-S. We introduce two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs. Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
@@
where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network. According to the phase analysis in section 4.2, the RC-S scheme stipulates that the number of VPs is equal to the number of intersection nodes at any given time. Therefore, we obtain:
$n_{VP}=n_{h}\cdot n_{v}$ (14)
@@
This factor represents the impact of the number of workers on the quantity of VPs that can be occupied. For each aisle, the newly generated VPs at the entrance could be occupied by loaded robots when its corresponding loading station is active and worker is engaged in tasks. The quantity of such VPs is a critical determinant of the maximum throughput. Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network. To ensure the balance of traffic flow and mitigate congestion at the aisle network’s periphery, workers are are evenly divided into four groups and positioned along the four edges of the site, as depicted in Figure 9. Four horizontal aisles and four vertical aisles are served by workers. We merge the area covered by these aisles and obtain the blue polygon.
Figure 9: The impact of number of workers on the VPs that can be occupied
For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of workers to the network’s maximum worker capacity.
@@
The proportion of the blue region, noted as the workforce factor $\kappa$, could be obtained by:
@@
We propose the following proposition regarding the relationship between the workforce factor and the number of occupied VPs:
PROPOSITION 1.
If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)min\{n_{v},n_{h}\}$, then an upper bound of the proportion of occupied VPs is $\kappa$;
The proof is presented in Online Appendix B. Proposition 1 provides with support for estimating the maximized network throughput. Next, we analyze the utilization rate of VPs using the RC-S scheme.
@@
This factor represents the utilization of VPs by the feasible paths generated by RC-S. We begin with examine the path reservation strategy within the framework of RC-S . In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time. To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the spatial utilization.
Figure 10: The ”pile up” process of spatiotemporal occupation under RC
Figure 11: The loss of capacity in different network
@@
$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\}$ (18)
@@
We now derive the average travel distance $\bar{l}$ of a sorting task, as the remaining part of the throughput estimation formula. Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance. Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns better with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions. Sorting demand are uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery. However, from the diagram, it can be observed that to ensure the outlets in area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$. We will next present the rationale for their calculation and the formulas. For a detailed derivation process, please refer to the appendix B.
@@
For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $min\{n_{h},n_{v}\}\geq 4$, (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\textit{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance could be obtained by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
@@
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$. Denote the value as $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 2 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B. Proposition 2 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and this critical point is not achieved when all loading stations are activated, due to workload balancing and uniformly distributed demands. When the proportion of active loading stations approaches 1, the growth rate of $n^{occupied}_{VP}$ will be slower than the growth rate of the service path length, resulting in a decrease in throughput. We will verify the formula through experiments in the next section.
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
Table 3: Results on 12+12 aisle network
@@
40 8.83 4.72 11.73 14.17 5.67 5.99
80 15.12 9.03 12.92 14.57 6.80 16.85
120 17.45 12.40 13.89 15.24 8.50 35.37
160 18.06 14.32 14.26 15.79 9.31 55.31
200 18.33 14.94 14.51 16.03 9.63 63.36
Table 4: Results on 20+20 aisle network
@@
50 7.68 3.94 18.85 21.75 15.53 11.61
100 14.91 7.79 19.33 21.95 15.79 29.94
200 26.82 15.07 20.77 22.54 19.76 99.61
300 32.32 22.58 21.48 23.33 29.27 226.76
400 33.66 24.37 23.41 23.92 37.40 363.45
Tables 3 and 4 report the system throughput, average service time and average run-time. The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles. The difference in algorithm performance is primarily due to the following improvements: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, while CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; (2) RC-S limits the total number and maximum density of robots and is free of conflicts, while high robot density around outlets in benchmark algorithm leads to significant resource expenditure on conflict resolution.
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail. We conduct 20 rounds of replication for each scenario. In each experiment, we set a warm-up period of one hour and then record the system performance for three hours of continuous operation.
@@
Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste. Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$. Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We notice that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation.
6.3 Efficiency underestimating in queueing models
Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (Buzacott and Shanthikumar,, 1993). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing. However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3. Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2. The results are shown in Table 5. The CQN model significantly underestimates system throughput in scenarios with few robots. It fails to capture variations in traffic congestion across different flow rates and overestimates road conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
Table 5: Error in throughput estimation
$n_{h}+n_{v}$ $20+20$ $24+24$
$n_{r}$ 50 100 200 300 50 100 200 300
Throughput (×$\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
CQN 5.98 11.87 22.98 31.47 4.96 9.88 19.52 28.61
RC-S 7.83 15.66 31.32 33.86 6.55 13.10 26.19 39.08
Relative error (%) CQN 29.03 26.95 19.65 4.34 31.82 30.92 27.73 21.61
RC-S 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
@@ 7 Layout Design Optimization
The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs. Considering transportation costs and supply stability, we analyze the following scenario: The warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$. The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
@@
$C_{o}=\sum_{\sigma\in\mathcal{S}}\theta^{\sigma}(P_{w}n_{w}^{\sigma}+P_{r}n_{r}^{\sigma})$ (31)
Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $P_{w}$ and $P_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot. The objective of the problem is modeled as follows:
$C_{d}(T^{\sigma},\theta^{\sigma},N_{o})=C_{f}+C_{o}$ (32)
@@
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma}\qquad\forall\sigma\in\mathcal{S}$ (33)
@@
$n_{w}^{\sigma}\leq n_{l}\qquad\forall\sigma\in\mathcal{S}$ (36)
$n_{h}=2k_{h}\ ,\quad k_{h}\in\mathcal{Z}_{+}$ (37)
$n_{v}=2k_{v}\ ,\quad k_{v}\in\mathcal{Z}_{+}$ (38)
$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+}^{2\left|\mathcal{S}\right|}$ (39)
Constraint (33) ensures that the system throughput meets predicted sorting requirements in different periods. Constraint (34) restricts the lower bound of the site size to accommodate an adequate number of outlets. Constraint (35) represents the maximum number of loading stations accommodated in the aisle network under RC-S. Constraint (36) limits the maximum number of workers to be stationed. Constraints (37)-(38) ensure the number of aisles even. The remaining constraint (39) requires the variables to be positive integers.
@@
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}$ (40)